OpenStudy (anonymous):

Use the conversion factor 1amu=6.66054 x 10^-24 to answer the following questions. a) 1.674 x 10^-24 g of neutrons is how many amu? b)what is the mass in grams of one lithium atom(lithium has an atomic weight of 6.94 amu) c)6.492 x 10^-23 g of potassium is how many amu?

OpenStudy (turingtest):

a)You mean\[1\text{amu}=1.66054\times10^{-24}\text g\]so the conversion from grams to amu will be\[1.674\times10^{-24}\text g(\frac{1\text{amu}}{1.66054\times10^{-24}\text g})\]

OpenStudy (anonymous):

so what about b and c

OpenStudy (anonymous):

can you help me with this one also please. Complete combustion of 8.00g of a hydrocarbon produced 25.6g of CO2 and 9.15g of H2O. What is the empirical formula of the hydrocarbon? I have to leave. I will be back. Thanks for A.

OpenStudy (turingtest):

b)\[\frac{6.94\text{amu}}{1\text{Li atom}}\cdot\frac{1.66054\times10^{-24}\text{g}}{1\text{amu}}\]check the units to see that this gives g/Li atom

OpenStudy (turingtest):

c)\[?\text g\times\frac{?\text{amu}}{?\text{g}}=?\text{amu}\]you should be able to figure out the numbers yourself after the last two problems

OpenStudy (anonymous):

ok can you help me with the other problem i posted please

OpenStudy (turingtest):

I'm looking at it, but I haven't done that stuff in a while.

OpenStudy (anonymous):

I already told him how to answer that problem, but it looks like he only wants answers anyways.

OpenStudy (turingtest):

thanks for the heads up on that

OpenStudy (anonymous):

what. where did you say?

OpenStudy (xishem):

In the original post you made with the exact same question: http://openstudy.com/study#/updates/4f302c9de4b0fc09381ec1b6